OpenStudy (scarlettfarra2000):

Does anyone know what to numbers can be added together to make 15 but also multiply together to make 164? Can someone help me I have no clue

OpenStudy (mahoganie.carson):

Hey i had to do something like this today for school use this calculator http://www.mathpapa.com/ Hope this help :)

OpenStudy (scarlettfarra2000):

Says there's no like terms

OpenStudy (mahoganie.carson):

Well 10 + 5 = 15 15 * 164 = 2460 :)

OpenStudy (mahoganie.carson):

@ScarlettFarra2000

OpenStudy (scarlettfarra2000):

What I'm trying to do is get \[x^2+15x+164= (x+?)(x-?) \] something like that but I don't know the numbers

OpenStudy (mahoganie.carson):

Okay i think the answer is this x^2 + 15x + 164

OpenStudy (mahoganie.carson):

@ScarlettFarra2000

OpenStudy (scarlettfarra2000):

Really?

OpenStudy (mahoganie.carson):

Yeah i did it on a calculator and the answer was that

OpenStudy (scarlettfarra2000):

Could you please explain I've been working on this for three days

OpenStudy (mahoganie.carson):

Okay i went to the mathpapa calculator right and i did this x^2 + 15x + 164 = and then the answer was x^2 + 15x + 164

OpenStudy (scarlettfarra2000):

See i got the same answer hold i'll screen shot the problem

OpenStudy (agent0smith):

There are no numbers that add to 15 and multiply to 164. The biggest possible product would be 8*7 which only gives 56. It doesn't factor.

OpenStudy (scarlettfarra2000):

Maybe I didn't do the problem right?

OpenStudy (agent0smith):

Area = length times width Area of pool+path = (14+2x)(16+2x) = 440 (they told you that)

OpenStudy (scarlettfarra2000):

I still can't found out where I did this wrong

OpenStudy (anonymous):

The answer is X=6 or X=-36 According to the problem in the file above it is : X^2+30x-216=0

OpenStudy (anonymous):

Sorry this is the right answer : (16+2x)*(14+2x)=440 224+32x+28x+4x^2=440 4x^2 + 60x - 216=0 X^2 + 15x - 54 =0 (X+18)(X-3)=0 X=-18 (not possible cause it's length or width) Or X=3

OpenStudy (scarlettfarra2000):

Thank you so much @hinamori you don't know how thankful I'm am

OpenStudy (anonymous):

My pleasure, and glad I could help :)